# A Case Study on Misinformation Detection

### v3

@@ 4.3 Users
Detection targets. Papers within this scope target source reputation, and (U.i) inspect account metadata, such bios, account age, and profile images; or distinguish suspicious accounts by (U.ii) a single user’s social behaviors, such as their comments on posts (n.b. this target is distinct from (N.i)). The security literature we review discusses trolls and bots deployed for astroturfing, misinformation campaigns, and IOs [361, 367, 368, 369, 370]. In the absence of rigorous definitions of these account types, however, actual detection targets are tautological: A troll or bot is an account that exhibits troll- or bot-like behavior, or that interacts with confirmed troll or bot accounts [361, 367].
@@
Datasets. All troll and bot account detection works we reviewed relied on published lists of “known” troll accounts for model training, but (2.iv) neglected to mention the heavily manual investigation required to produce these original lists [300, 385]. Researchers who compiled some of these account lists, including a set of several hundred Twitter accounts with possible links to a known Russian troll farm (the Internet Research Agency, or IRA) manually examined suspicious accounts and tweet contents in order to produce detailed account and content taxonomies; notably, these classifications required external intelligence about account activity that was not published alongside account lists [385]. Two other highly-cited troll lists compiled by the U.S. government, comprising several thousand suspicious Twitter and Facebook accounts, were curated through the use of proprietary non-public information [423].
@@
Model selection. Detection proceeds via classifier training on a list of “known” suspicious accounts and application of this classifier to a dataset of novel accounts [361, 364]. We note that, regardless of model choice, if classifier training data and feature selection reflect a heuristic about suspicious behavior, the resulting classifier will simply learn this heuristic: The methods we review can be used to detect accounts whose behaviors conform to heuristic assumptions, but cannot be used to surface novel malicious behaviors, and are not resistant to attacks or evasion [371]; we explore this further in our replication analysis of Saeed et al. (Section 5.2) [361]. A subset of methods at this scope and the network scope formulate detection as a (3.iv) graph cut or influence maximization problem, and describe approaches to identifying optimal cuts for isolating suspicious accounts [257, 101].
@@
Feature selection. Methods that define suspicious accounts by intrinsic properties of these accounts (e.g., user handles and profile images) target the (4.iii) semantics of this account metadata, and detect evidence of manipulation in (e.g.) image metadata and bios; or text outputs, such as posts and links [5, 384]. Methods that define suspicious accounts by account activity target (4.ii) networked behaviors, such as liking and resharing statistics [92]. Feature sets for some methods in the first category include demographic data for users, such as inferred political party affiliation or race [254] (these nth order assumptions are dangerous to make [100]; see our discussion about proxy signals for detection, in Section 4.2.
@@
Evaluation. We observe accuracy scores above 80% for (5.ii) confirmatory detection of like accounts for all methods that reference a seed list of known trolls [361, 66, 366]; de novo detection of behaviors not represented within training data is not possible, by the self-admission of 10% of authors within this scope [361, 359]. As we discuss further in the next subsection (Networks), the increasingly hybrid nature of IOs requires more nuanced taxonomies for classification: extent of coordination, rather than existence, might be a more appropriate measure of possible manipulation. Formerly, Botometer did this by assigning a “bot-ness” score to individual Twitter accounts that gauged extent of automation [422]. Though currently defunct post-Twitter API shutdown, Botometer performance had already degraded significantly on Twitter data from 2020, per Rauchfleisch et al. [372]. [^12] In a blog post on bot detection from 2021, Twitter “debunked” four common heuristics commonly used to identify bot accounts, including several detected by methods in our corpus, and described a “forensic team of investigators” who manually verify bot-ness of suspect accounts [398]. Some authors of bot detection methods acknowledge that their approaches are (5.iii) trivially easy to evade if account holders 1) avoid interacting with known suspicious accounts or 2) vary their account identity and posting semantics [373, 367].
Takeaways: 1 Targets are frequently tautological. 2 Hand-annotated training datasets are the result of intensive fact-finding on the part of human researchers, and often require information that is not publicly available. 3 Classifiers can only detect accounts resembling those in seed lists. 4 Features that attempt to infer user credibility from demographic information risk reinforcing existing biases. 5 Current methods cannot detect novel malicious behaviors.
4.4 Networks
Within the security literature, a growing awareness of hybrid networks, which employ a combination of automated and manual approaches to disseminate content, has encouraged a turn toward network-based bot detection methods, and away from detection of individual accounts [367, 374]. We observe a parallel turn toward network-based methods in AI, ML, and NLP venues as a result of growing recognition of overfitting and generalizability issues in purely text-based detection methods (see Articles) [302]. The common assumption, across disciplines, is that coordinated networks leave more detectable evidence of manipulation than do individual accounts, and that these footprints should be identifiable regardless of attack type or rumoring topic [302, 378]. Network-scoped methods, including relevant work in the security literature, form 20% of our corpus.
@@ 4.4 Networks
Detection targets. All methods at this scope identify patterns of user interaction and content propagation as targets; these methods associate virality with the existence of rumoring narratives [6, 228, 41] and temporally anomalous activity with evidence of coordination [359, 367, 364]. The corresponding targets for these approaches are (N.i) propagation patterns across social graph topologies and (N.ii) temporal records of user-user interactions. Within non-security misinformation literature, we note that virality assumptions disallow early detection of misinformation [315, 253]. Similarly, within the security literature, anomalous patterns of account registration and user interaction serve as proxies for the presence of Sybils and botnets [375, 373]; early detection requires that authors formulate a priori assumptions about the nature of these patterns.
@@
Datasets. We conducted an author outreach survey for works within this scope in an attempt to locate hard-to-find social media datasets. We found that (2.v) accessibility issues were exacerbated by the shutdown of the Twitter API [308]. In total, we attempted to locate datasets and code for 50 different papers (see Section 5 for our methodology). Thirty-six (72%) of these analyzed tweet corpora, and 42 (84%) of these targeted social media users and posting contents. We were able to independently source complete methods or data for fourteen (28%) of these. Of the 27 authors we eventually contacted about providing partial or dehydrated datasets (as we were unable to locate these datasets ourselves), nine responded; six of those authors were able to provide method code or partial datasets.
@@
Model selection. The network-scoped methods we review formulate detection as 1) a structured content classification problem [370, 388], and/or 2) a clustering problem on social graphs [391, 157]. In the former case, authors employ an assortment of (3.ii, 3.iii) supervised and unsupervised models to detect suspicious language across multiple accounts. In the latter case, authors use (3.vi) Louvain or Kmeans clustering or K-nearest neighbors to detect neighborhoods of suspicious accounts, as determined by user-user interactions. Though methods in the latter category advertise themselves as content-agnostic, we note that published methods [40, 30] access datasets of social media posts that were already sorted by rumoring topic or event [189, 202].
@@
Feature selection. 55% of papers within this scope make normative theoretical assumptions about user behaviors: In keeping with an epidemiological model [^13] Some academics have argued that the disease metaphor for misinformation promotes an overly simplistic model of information spread and opinion formation [399, 400]. of misinformation spread [132, 328], Nguyen et al. assume a homogeneous population of newsreaders, with (4.ii) identical probabilities of “infection” and reinfection [205]. Similarly, in the security literature, techniques for detecting bots and Sybils identify behaviors that align with heuristics determined a priori by researchers: These methods assume, for instance, that Sybils will form well-connected neighborhoods [375], or (seemingly contradictorily) that compromised Sybils will refrain from connecting with additional Sybils, to avoid detection [373]. [^14] In fact, the landmark paper by Douceur that characterized Sybil attacks stated that such attacks cannot be prevented unless special assumptions are made about account behaviors [401].
@@
Evaluation. Ferrara et al. [379] called attention to the false positive rate problem in botnet detection in 2016, noting that classifiers for bot detection only work well in instances where there is a clear-cut distinction between bot and non-bot accounts. This distinction is becoming increasingly blurred, however. Sophisticated network-based attacks try to engage non-bot accounts in organic interactions with bot accounts (e.g., astroturfing attacks) [367, 359], (5.v) rendering even positive detection results insufficient or meaningless: coordinated activity need not be inauthentic, and inauthentic activity need not be malicious.
Takeaways: 1 Pattern- and virality-based detection approaches disallow early detection of rumors. 2 Current social media data is difficult to obtain. 3 “Topic-agnostic” classifier design occurs downstream of topic-aware dataset design. 4 Epidemiological models of information spread make strong assumptions about opinion formation and user behaviors; feature sets reflect these a priori notions. 5 Though network-scoped methods are less susceptible to content-based dependencies than are content-aware methods, they cannot infer intent or authenticity of the behaviors they detect.
4.5 Websites
Methods within this scope apply credibility, factuality, or (political) bias scores to whole news sites; authors claim that site-wide labels can be used to quickly infer the quality of individual news articles produced by these sites [1, 316, 302, 109]. As with article-scoped methods, we noted intervention fit issues at the whole website scope. Asr et al. found that whole-source labels were insufficient proxies for the truthfulness of single news articles, and elided subject-specific variations in reporting quality [287]. (We discuss dataset distribution in greater depth in Datasets.)
1
Detection targets. In 50% of works that we review in this scope, authors reduce the task of whole-site credibility labeling to a significantly smaller, unimodal classification task: Chen et al. [109] (W.i) detect suspicious domain semantics (in essence, a text classification task on URLs); Ribeiro et al. [254] (W.ii) infer site bias from site visitor demographics; Castillo et al. [253] (W.iii) detect suspicious ad interfaces and markup features; Baly et al. [43, 316] and Hounsel et al. [1] present methods incorporating (W.iv) infrastructural features, though the overall performance of the method of Baly et al. is strongly determined by performance on the text classification task alone (thus, in practice, the method closely resembles the article-scoped detection methods we examine, and is susceptible to the same dependencies that we observe in that scope). We demonstrate this via an ablation analysis in our replication study of their method (see Section 5.3).
2
Datasets. For both training and testing, all methods scoped to whole website detection rely on published lists of websites with accompanying credibility scores [43, 316, 1, 302]; common reference sites include Media Bias/Fact Check, Snopes, and FactCheck.org [311, 298, 402]. As discussed in Section 4.1, however, these references do not have uniform taxonomies for classifying site credibility. Additionally, pre-labeled lists are 1) (2.i) biased towards older, more visible real news and fake news outlets [298], 2) (2.iii) are restricted to specific information domains [43], or 3) (2.i) include inactive websites within their labeled datasets [311]. [^15] The median lifespan of a set of 283 misinformation news sites is 4 years, per a survey conducted by Chalkiadakis et al. in 2021 [114]. We note papers that perform website infrastructure analysis on contemporaneous snapshots (circa 2019) of websites in their corpora, even though the text-based features for the same analysis were drawn from datasets published in 2016 and 2017 [302]. This constitutes a serious source of (2.ii) temporal leakage. No works within this scope discuss approaches to accounting for uneven distributions in training data, or how they might account for shifts in baseline distributions during the lifecycle of an active website. Hounsel et al., for instance, train their classifier on a reference list in which 34% of misinformation training set sites were active and all websites in the real news training set were active, possibly resulting in overfitting to features specific to those inactive websites [1].
3
Model selection. Four of the seven methods we reviewed within this scope employed (3.ii) SVM classifiers; in two of those cases, SVM outperformed other, more complex unsupervised models [287, 302]. These results accord with our earlier observation, in Section 4.2, that SVM classifiers are comparatively fairly performant on two-way classification tasks.
4
Feature selection. Works within this scope employ multi-modal feature sets comprising a mix of (4.i) textual, (4.ii) network-based, and (4.iv) infrastructural signals: for instance, Baly et al. [316] consider network traffic, URL semantics, and site contents; and Hounsel et al. [1] consider TLS/SSL certificates, web hosting configurations, and domain registrations. None of the detection methods we reviewed discusses the computational costs of deploying their methods at scale. Though all works discuss their feature selection process (via leave-one-out and use-one-only evaluations), none describes a process for normalizing or weighting features according to dataset distribution or detection setting needs.
5
@@ 4.5 Websites
Evaluation. Methods within this scope that propose to perform whole-site labeling from analysis of a selection of news articles or infrastructural features are susceptible to distributional imbalances (e.g., between news verticals represented in an article corpus). Baly et al. train their model on (5.ii) political news websites only, and their credibility labels are strongly correlated with political bias scores. Hounsel et al. perform (5.i) testing in real time—one of the few studies we reviewed, and one of two studies at this scope, that did so [1, 109]. Both of these works report significant performance dropoffs between experimental and real-time tests, most likely as a result of distributional differences between real-world and experimental datasets (in practice, most websites do not host news-related content at all) [1, 109].
@@
Takeaways: 1 Methods claiming to classify news sites generally reduce this task to simpler, unimodal ones, such as URL classification. 2 The works we review do not consider shifts in feature distribution over time. 3 We note that SVM classifiers perform well on two-way classification tasks, and even outperform more sophisticated unsupervised models. 4 Feature normalization largely undiscussed. 5 Authors who conduct testing in separate real-time settings report significant performance dropoffs with respect to experimental results.
@@ 5 REPLICATION STUDIES
We choose three distinct targets and scopes to replicate issues identified in our literature review. Our targets are 1) (A.i) syntax-based text features; 2) (U.ii) user behaviors; and 3) (W.iv) multimodal whole-website features, including hosted content and infrastructure. These works are highly representative of their respective information scopes: Nasir et al. [141] (Section 5.1) train a neural network on corpora of true and misinformative news articles; Saeed et al. [361] (Section 5.2) use published lists of trolls to infer the presence of other troll accounts on Reddit; and Baly et al. [43] (Section 5.3) examine a multimodal feature set comprising infrastructure, content, and network-based features in order to infer whole-site credibility. For each work, we evaluate replicability and generalizability (RQ3, RQ4). Where possible, we 1) replicate reported results; 2) inspect datasets for potential dependencies; 3) perform ablation analyses to understand individual feature performance; and 4) test on current data.
@@
Ethical Considerations. Datasets used for this work were already publicly available or were obtained with permission from study authors. Our outreach study was approved by our institutional review board.
@@ 5.1 Detection of suspicious language
In view of rampant data dependency issues identified in our survey of article-scoped literature, we reproduced results from a representative study published in 2021 by Nasir et al [141]. The authors propose a neural net-based approach to the classification of news articles. The method employs a hybrid deep learning model that combines convolutional and recurrent neural networks for the classification of real and fake news. The authors report results for tests on two datasets: the ISOT dataset (45,000 news stories, equally distributed across true and false categories, as labeled by PolitiFact) and the FA-KES dataset (804 news stories about the Syrian war, 426 true and 376 false) [312, 313]. We were able to replicate original paper results and run the method on updated datasets of news articles. Motivated by the prevalence of methods trained on few- or single-source datasets in the article scope, we test possible dependencies related to journalistic house style, as all true articles in the Nasir et al. training corpus were sourced from Reuters.
House style as a confounder. To investigate house style as a possible confounder for misinformation detection, we excerpted 100 articles from both Reuters and The New York Times, two news outlets with distinctive (and different) reporting styles. We randomly selected these articles from both outlets’ RSS feeds in May 2023. We sourced articles for both corpora from the following verticals: U.S. and world politics, economics, science, and entertainment. Excerpt lengths ranged from 100 to 300 words. These corpora, each comprising 100 true news stories, are labeled “Reuters-original” and “NYTimes-original” in Table II.
@@
TABLE II: Replication analysis of Nasir et al. (2021). We tested the method of Nasir et al. on both datasets discussed in the original paper, and on novel datasets from Reuters and The New York Times [141].
We then selectively edited 50% of the news articles within each corpus. We changed proper nouns, negated verbs, and altered reported statistics so that the factual content of these articles was no longer accurate but house style and tone were preserved. We call these altered text corpora “Reuters-modified” and “NYTimes-modified” in Table II. The classifier had 0.727 accuracy on Reuters-original and 0.673 accuracy on NYTimes-original; this difference is not significant (p > 0.05). Additionally, the classifier had about 0.50 (random) accuracy on both modified datasets. The classifier’s false positive and false negative rates on modified and original Reuters and NYTimes corpora tell a more interesting story, however: Overwhelmingly, false NYTimes articles were classified as true (FPR = 0.02 and 0.735 for Reuters-modified and NYTimes-modified, respectively), while all true Reuters articles were classified as false (FNR = 1 and 0.286 for Reuters-modified and NYTimes-modified, respectively). These results indicate that the classifier was significantly more conservative in its assignment of true labels for the Reuters dataset than it was for the NYTimes dataset; additionally, within each modified corpus, the classifier did not effectively differentiate between true and untrue news articles. We did not investigate the stylistic features of each corpus that might have produced (anti-)conservative label distributions; for this work, however, it is sufficient to note that classifier performance in an adversarial setting was no better than random, and that house style appears to have a significant impact on classifier sensitivity to misinformative texts.

### v4

@@ 4.4 Networks
Detection targets. All methods at this scope identify patterns of user interaction and content propagation as targets; these methods associate virality with the existence of rumoring narratives [vosoughi-twitter-2018, tambuscio_fact-checking_2015, shu_beyond_2019] and temporally anomalous activity with evidence of coordination [stone2009botnet, cresci_bots_2020, lukito2020coordinating]. The corresponding targets for these approaches are (N.i) propagation patterns across social graph topologies and (N.ii) temporal records of user-user interactions. Within non-security misinformation literature, we note that virality assumptions disallow early detection of misinformation [monti2019fake, castillo_information_2011]. Similarly, within the security literature, anomalous patterns of account registration and user interaction serve as proxies for the presence of Sybils and botnets [yuan2019detecting, danezis2009sybilinfer]; early detection requires that authors formulate a priori assumptions about the nature of these patterns.
@@
Datasets. We conducted an author outreach survey for works within this scope in an attempt to locate hard-to-find social media datasets. We found that (2.v) accessibility issues were exacerbated by the shutdown of the Twitter API [twitter]. In total, we attempted to locate datasets and code for 50 different papers (see Section 5 for our methodology). Thirty-six (72%) of these analyzed tweet corpora, and 42 (84%) of these targeted social media users and posting contents. We were able to independently source complete methods or data for fourteen (28%) of these. Of the 27 authors we eventually contacted about providing partial or dehydrated datasets, nine responded; six of those authors were able to provide method code or partial datasets.
@@
Model selection. The network-scoped methods we review formulate detection as 1) a structured content classification problem [alizadeh2020content, ratkiewicz2011detecting], and/or 2) a clustering problem on social graphs [assenmacher2020two, sharma_combating_2019]. In the former case, authors employ an assortment of (3.ii, 3.iii) supervised and unsupervised models to detect suspicious language across multiple accounts. In the latter case, authors use (3.vi) Louvain or K-means clustering or K-nearest neighbors to detect neighborhoods of suspicious accounts, as determined by user-user interactions. Though methods in the latter category advertise themselves as content-agnostic, we note that published methods [liu_early_nodate, ma_detect_2017] access datasets of social media posts that were already sorted by rumoring topic or event [ma_rnns_2016, liu_realtime_2015].
@@
Feature selection. 55% of papers within this scope make normative theoretical assumptions about user behaviors: In keeping with an epidemiological model [^11] Some academics have argued that the disease metaphor for misinformation promotes an overly simplistic model of information spread and opinion formation [williams_2023, tay2024thinking]. of misinformation spread [jin_epidemiological_2013, vanderlinden_2022], Nguyen et al. assume a homogeneous population of newsreaders, with (4.ii) identical probabilities of “infection” and reinfection [nguyen_containment_2012]. Similarly, in the security literature, techniques for detecting bots and Sybils identify behaviors that align with heuristics determined a priori by researchers: These methods assume, for instance, that Sybils will form well-connected neighborhoods [yuan2019detecting], or (seemingly contradictorily) that compromised Sybils will refrain from connecting with additional Sybils, to avoid detection [danezis2009sybilinfer]. [^12] In fact, the landmark paper by Douceur that characterized Sybil attacks stated that such attacks cannot be prevented unless special assumptions are made about account behaviors [douceur2002sybil].
@@
Evaluation. Ferrara et al. [ferrara2016rise] called attention to the false positive rate problem in botnet detection in 2016, noting that classifiers for bot detection only work well in instances where there is a clear-cut distinction between bot and non-bot accounts. This distinction is becoming increasingly blurred, however. Sophisticated network-based attacks try to engage non-bot accounts in organic interactions with bot accounts (e.g., astroturfing attacks) [cresci_bots_2020, stone2009botnet], (5.v) rendering even positive detection results insufficient or meaningless: coordinated activity need not be inauthentic, and inauthentic activity need not be malicious.
**Takeaways: 1 Pattern- and virality-based detection approaches disallow early detection of rumors. 2 Current social media data is difficult to obtain. 3 “Topic-agnostic” classifier design occurs downstream of topic-aware dataset design. 4 Epidemiological models of information spread make strong assumptions about opinion formation and user behaviors; feature sets reflect these a priori notions. 5 Though network-scoped methods are less susceptible to content-based dependencies than are content-aware methods, they cannot infer intent or authenticity of the behaviors they detect.
4.5 Websites
Methods within this scope apply credibility, factuality, or (political) bias scores to whole news sites; authors claim that site-wide labels can be used to quickly infer the quality of individual news articles produced by these sites [hounsel-infra-2020, baly-etal-2020-written, castelo_agnostic_2019, chen_proactive_2020]. As with article-scoped methods, we noted intervention fit issues at the whole website scope. Asr et al. found that whole-source labels were insufficient proxies for the truthfulness of single news articles, and elided subject-specific variations in reporting quality [torabi_asr_data_2018]. (We discuss dataset distribution in greater depth in Datasets.)
@@ 4.5 Websites
Detection targets. In 50% of works that we review in this scope, authors reduce the task of whole-site credibility labeling to a significantly smaller, unimodal classification task: Chen et al. [chen_proactive_2020] (W.i) detect suspicious domain semantics (in essence, a text classification task on URLs); Ribeiro et al. [ribeiro_media_2018] (W.ii) infer site bias from site visitor demographics; Castillo et al. [castillo_information_2011] (W.iii) detect suspicious ad interfaces and markup features; Baly et al. [baly_predicting_nodate, baly-etal-2020-written] and Hounsel et al. [hounsel-infra-2020] present methods incorporating (W.iv) infrastructural features, though the overall performance of the method of Baly et al. is strongly determined by performance on the text classification task alone (thus, in practice, the method closely resembles the article-scoped detection methods we examine, and is susceptible to the same dependencies that we observe in that scope). We demonstrate this via an ablation analysis in our replication study of their method (see Section 5.3).
@@
Datasets. For both training and testing, all methods scoped to whole website detection rely on published lists of websites with accompanying credibility scores [baly_predicting_nodate, baly-etal-2020-written, hounsel-infra-2020, castelo_agnostic_2019]; common reference sites include Media Bias/Fact Check, Snopes, and FactCheck.org [mediabiasfactheck, snopes, FactCheck]. As discussed in Section 4.1, however, these references do not have uniform taxonomies for classifying site credibility. Additionally, pre-labeled lists are 1) (2.i) biased towards older, more visible real news and fake news outlets [snopes], 2) (2.iii) are restricted to specific information domains [baly_predicting_nodate], or 3) (2.i) include inactive websites within their labeled datasets [mediabiasfactheck]. [^13] The median lifespan of a set of 283 misinformation news sites is 4 years, per a survey conducted by Chalkiadakis et al. in 2021 [chalkiadakis_rise_2021]. We note papers that perform website infrastructure analysis on contemporaneous snapshots (circa 2019) of websites in their corpora, even though the text-based features for the same analysis were drawn from datasets published in 2016 and 2017 [castelo_agnostic_2019]. This constitutes a serious source of (2.ii) temporal leakage. No works within this scope discuss approaches to accounting for uneven distributions in training data, or how they might account for shifts in baseline distributions during the lifecycle of an active website. Hounsel et al., for instance, train their classifier on a reference list in which 34% of misinformation training set sites were active and all websites in the real news training set were active, possibly resulting in overfitting to features specific to those inactive websites [hounsel-infra-2020].
@@
Model selection. Four of the seven methods we reviewed within this scope employed (3.ii) SVM classifiers; in two of those cases, SVM outperformed other, more complex unsupervised models [torabi_asr_data_2018, castelo_agnostic_2019]. These results accord with our earlier observation, in Section 4.2, that SVM classifiers are comparatively performant on two-way classification tasks.
@@
Feature selection. Works within this scope employ multi-modal feature sets comprising a mix of (4.i) textual, (4.ii) network-based, and (4.iv) infrastructural signals: for instance, Baly et al. [baly-etal-2020-written] consider network traffic, URL semantics, and site contents; and Hounsel et al. [hounsel-infra-2020] consider TLS/SSL certificates, web hosting configurations, and domain registrations. None of the detection methods we reviewed discusses the computational costs of deploying their methods at scale. Though all works discuss their feature selection process (via leave-one-out and use-one-only evaluations), none describes a process for normalizing or weighting features according to dataset distribution or detection setting needs.
@@
Evaluation. Methods within this scope that propose to perform whole-site labeling from analysis of a selection of news articles or infrastructural features are susceptible to distributional imbalances (e.g., between news verticals represented in an article corpus). Baly et al. train their model on (5.ii) political news websites only, and their credibility labels are strongly correlated with political bias scores. Hounsel et al. perform (5.i) testing in real time—one of the few studies we reviewed, and one of two studies at this scope, that did so [hounsel-infra-2020, chen_proactive_2020]. Both of these works report significant performance dropoffs between experimental and real-time tests, most likely as a result of distributional differences between real-world and experimental datasets (in practice, most websites do not host news-related content at all) [hounsel-infra-2020, chen_proactive_2020].
@@
**Takeaways: 1 Methods claiming to classify news sites generally reduce this task to simpler, unimodal ones, such as URL classification. 2 The works we review do not consider shifts in feature distribution over time. 3 We note that SVM classifiers perform well on two-way classification tasks, and even outperform more sophisticated unsupervised models. 4 Feature normalization largely undiscussed. 5 Authors who conduct testing in separate real-time settings report significant performance dropoffs with respect to experimental results.
@@ 5 REPLICATION STUDIES
We choose three distinct targets and scopes to replicate issues identified in our literature review. Our targets are 1) (A.i) syntax-based text features; 2) (U.ii) user behaviors; and 3) (W.iv) multimodal whole-website features, including hosted content and infrastructure. These works are highly representative of their respective information scopes: Nasir et al. [nasir_fake_2021] (Section 5.1) train a neural network on corpora of true and misinformative news articles; Saeed et al. [saeed2022trollmagnifier] (Section 5.2) use published lists of trolls to infer the presence of other troll accounts on Reddit; and Baly et al. [baly_predicting_nodate] (Section 5.3) examine a multimodal feature set comprising infrastructure, content, and network-based features in order to infer whole-site credibility. For each work, we evaluate replicability and generalizability (RQ3, RQ4). Where possible, we 1) replicate reported results; 2) inspect datasets for potential dependencies; 3) perform ablation analyses to understand individual feature performance; and 4) test on current data.
@@ 5.1 Detection of suspicious language
In view of rampant data dependency issues identified in our survey of article-scoped literature, we reproduced results from a representative study published in 2021 by Nasir et al [nasir_fake_2021]. The authors propose a neural net-based approach to the classification of news articles. The method employs a hybrid deep learning model that combines convolutional and recurrent neural networks for the classification of real and fake news. The authors report results for tests on two datasets: the ISOT dataset (45,000 news stories, equally distributed across true and false categories, as labeled by PolitiFact) and the FA-KES dataset (804 news stories about the Syrian war, 426 true and 376 false) [isot, fa_kes]. We were able to replicate original paper results and run the method on updated datasets of news articles. Motivated by the prevalence of methods trained on few- or single-source datasets in the article scope, we test possible dependencies related to journalistic house style, as all true articles in the training corpus were sourced from Reuters.
House style as a confounder. To investigate house style as a possible confounder for misinformation detection, we excerpted 100 articles from both Reuters and The New York Times, two news outlets with distinctive (and different) reporting styles. We randomly selected these articles from both outlets’ RSS feeds in May 2023. We sourced articles for both corpora from the following verticals: U.S. and world politics, economics, science, and entertainment. Excerpt lengths ranged from 100 to 300 words. These corpora, each comprising 100 true news stories, are labeled “Reuters-original” and “NYTimes-original” in Table 3.
@@
Table 3: Replication analysis of Nasir et al. (2021). We tested the method of Nasir et al. on both datasets discussed in the original paper, and on novel datasets from Reuters and The New York Times [nasir_fake_2021].
We then selectively edited 50% of the news articles within each corpus. We changed proper nouns, negated verbs, and altered reported statistics so that the factual content of these articles was no longer accurate but house style and tone were preserved. We call these altered text corpora “Reuters-modified” and “NYTimes-modified” in Table 3. The classifier had 0.727 accuracy on Reuters-original and 0.673 accuracy on NYTimes-original; this difference is not significant (p > 0.05). Additionally, the classifier had about 0.50 (random) accuracy on both modified datasets. The classifier’s false positive and false negative rates on modified and original Reuters and NYTimes corpora tell a more interesting story, however: Overwhelmingly, false NYTimes articles were classified as true (FPR = 0.02 and 0.735 for Reuters-modified and NYTimes-modified, respectively), while all true Reuters articles were classified as false (FNR = 1 and 0.286 for Reuters-modified and NYTimes-modified, respectively). These results indicate that the classifier was significantly more conservative in its assignment of true labels for the Reuters dataset than it was for the NYTimes dataset; additionally, within each modified corpus, the classifier did not effectively differentiate between true and untrue news articles. We note that classifier performance in an adversarial setting was no better than random, and that house style appears to have a significant impact on classifier sensitivity to misinformative texts.